a rectangular box has a volume of 480 cubic inches. the height of the box is 5 inches. the ratio of the length of the box to the width of thbox is 3 to 2. What is the measure of the width of the box?
Answer by mananth(16949)

L*W=96
L/w= 3/2
L=3W/2
..
Plug value of L
3W/2 *W= 96
3w^2/2=96
3W^2= 192
W^2= 192/3
W^2= 64
take sqrt
W= 8
plug value of W
L=3W/2
L=3*8/2
L=12
L=12, W= 8 , H =5
CHECK
12*8*5= 480 cu.inches